In manufacturing its product, City Manufacturing’s recipe requires 3 ounces of Material X at a standard cost of $0.20 per ounce. During January 2010, City bought and used 2,000 ounces of Material X at a cost of $0.18 per ounce and made 700 units of product. The company has estimated that it will produce 750 units of product per month.What is City's material quantity variance?
Answer #1

Materials Quantity Variance = (SQ-AQ) * SP

= (700*3 - 2000) * 0.20

= (2100-2000) * 0.20

= 20 Favourable